# Profinite rigidity witnessed by Dehn fillings of cusped hyperbolic 3-manifolds

### v2

@@ 6.4. The Berge manifold
Proposition 6.13 shows that there are three different slopes on one boundary component of $S^{3}\setminus\mathring{N}(L_{A})$ such that the corresponding Dehn filling is homeomorphic to $D^{2}\times S^{1}$. According to Theorem A, there exist three different slopes $c_{1},c_{2},c_{3}$ on one boundary component $\partial_{0}N$ of $N$ such that $\widehat{\pi_{1}N_{c_{1}}}\cong\widehat{\pi_{1}N_{c_{2}}}\cong\widehat{\pi_{1}N_{c_{3}}}\cong\widehat{\pi_{1}(D^{2}\times S^{1})}\cong\widehat{\mathbb{Z}}$. Note that the group $\mathbb{Z}$ is profinitely rigid among all finitely generated, residually finite groups. So according to Proposition 2.1, we derive that $\pi_{1}N_{c_{1}}\cong\pi_{1}N_{c_{2}}\cong\pi_{1}N_{c_{3}}\cong\mathbb{Z}$. For each $i=1,2,3$, $N_{c_{i}}$ is a compact orientable 3-manifold whose boundary is a torus, and the only such manifold with infinite cyclic fundamental group is $D^{2}\times S^{1}$. Thus, $N$ admits three different Dehn fillings on $\partial_{0}N$ yielding $D^{2}\times S^{1}$, and Proposition 6.13 implies that $N\cong S^{3}\setminus\mathring{N}(L_{A})$. ∎
@@ 6.5. Detecting Klein bottles and Dehn surgeries on $\mathcal{L}$ and $L_{B}$
Recall that $\mathcal{N}$ is the orientable $I$-bundle over Klein bottle, and $M_{1}$ is the complement of the Whitehead link $\mathcal{W}$ in $S^{3}$. In the following context, we denote $M_{3_{1}}$ as the complement of the trefoil in $S^{3}$, i.e. $M_{3_{1}}$ is a Seifert fibered space $(D^{2};(2,1),(3,1))$.
@@
According to (6.15), $L_{B}(0,r,0)\cong L_{B}(0,0,r)$ and $L_{B}(0,r,4)\cong L_{B}(0,4,r)$. Thus, they can be viewed as Dehn fillings on $L_{B}(0,0)$ and $L_{B}(0,4)$. According to (1), $L_{B}(0,0,r)$ and $L_{B}(0,4,r)$ can be decomposed by an embedded torus into $\mathcal{N}\cup M^{\prime}$, where $M^{\prime}$ is a Dehn filling of $M_{1}$ on one of its boundary components $\partial_{1}M_{1}$. The Dehn fillings of $M_{1}$ on $\partial_{1}M_{1}$ have been classified by [37, Table A.1]. In fact, there is exactly one slope yielding the solid torus $D^{2}\times S^{1}$, while the other slopes yield either a hyperbolic manifold, a Seifert fibered space with hyperbolic base orbifold, or a graph manifold with non-trivial JSJ-decomposition. In other words, there is only one element $r_{0}\in\mathbb{Q}\cup\{\infty\}$ such that $L_{B}(0,0,r_{0})\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$, which is indeed non-aspherical; while for any $r\neq r_{0}\in\mathbb{Q}\cup\{\infty\}$, $L_{B}(0,0,r)\cong\mathcal{N}\cup_{T^{2}}M^{\prime}$ is aspherical. We claim that $r_{0}=\infty$. In fact, by (6.15), $L_{B}(0,0,\infty)\cong L_{B}(0,\infty,0)$ is a Dehn filling on $L_{B}(0,\infty)$. [32, Lemma 2.3 (1)] showed that $L_{B}(0,\infty)\cong\mathcal{N}$, and thus $L_{B}(0,\infty,0)\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$. Similarly, there is only one element $r_{4}\in\mathbb{Q}\cup\{\infty\}$ such that $L_{B}(0,4,r_{4})\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$ is spherical, while for any $r\neq r_{4}\in\mathbb{Q}\cup\{\infty\}$, $L_{B}(0,4,r_{4})\cong\mathcal{N}\cup_{T^{2}}M^{\prime}$ is aspherical. One can similarly show that $r_{4}=\infty$ by [32, Lemma 2.3 (1)]. Thus, this proves (2).
(3) follows from [37, Table 3]. In fact, the link $\mathcal{L}$ is obtained by a $-\frac{1}{2}$-surgery on one component of the three-chain link Figure 10 in $S^{3}$. Thus, for any $r,s\in\mathbb{Q}\cup\{\infty\}$, $\mathcal{L}(r,s)$ can be viewed as $(-\frac{1}{2},r-2,s-2)$-surgery on three components of the three-chain link, for which [37] presents a complete classification of exceptional Dehn surgeries.
@@
    Let $c_{1},c_{2}\in\text{slp}(T)$ and $c_{1}^{\prime},c_{2}^{\prime}\in\text{slp}(T^{\prime})$ such that $\Delta(c_{1},c_{2})=\Delta(c_{1}^{\prime},c_{2}^{\prime})=4$.
    Then there exist exactly two (equivalence classes of) isomorphisms $\psi_{1},\psi_{2}\in\mathrm{Isom}_{\mathbb{Z}}(\pi_{1}(T),\pi_{1}(T^{\prime}))/(\psi\sim-\psi)\cong\mathrm{PGL}(2,\mathbb{Z})$ such that $\overline{\psi}_{i}(\{c_{1},c_{2}\})=\{c_{1}^{\prime},c_{2}^{\prime}\}$.
@@
We can now list all possible elements in $E$. Suppose $c=[p\mu+q\lambda]\in E$, where $(p,q)$ is a pair of coprime integers. Without loss of generality, we may assume $p\geq 0$. Then $\Delta(c_{1},c)=\left|q\right|\leq 4$, and $\Delta(c_{2},c)=\left|4p-q\right|\leq 4$. It is easy to verify that $(p,q)\in\{(0,\pm 1),(1,0),(1,1)(1,2),(1,3),(1,4)\}$. Thus, $E\subseteq\{[\mu],[\mu+\lambda],[\mu+2\lambda],[\mu+3\lambda],[\mu+4\lambda],[\lambda]\}$, and it is easy to verify that $[\mu]$ and $[\mu+4\lambda]$ is the unique pair realizing geometric intersection number $4$.
@@
Therefore, $H_{1}(\mathcal{L}(n-\frac{1}{2});\mathbb{Z})\not\cong H_{1}(L_{B}(0,r);\mathbb{Z})$, which implies $\widehat{\pi_{1}\mathcal{L}(n-\frac{1}{2})}\not\cong\widehat{\pi_{1}L_{B}(0,r)}$.
Now we consider the family $\mathcal{L}(n-\frac{1}{2})$ ($n\in\mathbb{Z}$). Let $M=\mathcal{L}(m-\frac{1}{2})$ and $N=\mathcal{L}(n-\frac{1}{2})$, and suppose $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then according to Theorem A, there exists an isomorphism $\psi:\pi_{1}\partial M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial N$ such that $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ for all $c\in\text{slp}(\partial M)$. In particular, $\overline{\psi}(\mathcal{E}_{k}(M,\partial M))=\mathcal{E}_{k}(N,\partial N)$ according to Proposition 5.10. We follow the ${}\mathbb{Q}\cup\{\infty\}$-parametrization for $\text{slp}(\partial M)$ and $\text{slp}(\partial N)$ as above. Recall that $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{k}(M,\partial M)\}=\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{k}(N,\partial N)\}=4$ by Lemma 6.21. Thus, according to Corollary 6.19 (3) and Lemma 6.22 (1), $\{-2,2\}$ is the unique pair of elements in $\mathcal{E}_{k}(M,\partial M)$, as well as in $\mathcal{E}_{k}(N,\partial N)$, that realizes the maximal geometric intersection number $4$. Since $\overline{\psi}:\mathcal{E}_{k}(M,\partial M)\xrightarrow{1:1}\mathcal{E}_{k}(N,\partial N)$ preserves the geometric intersection number by Lemma 6.2, we derive that $\overline{\psi}(\{-2,2\})=\{-2,2\}$. According to Lemma 6.22 (2), there are only two possible candidates for $\psi$ in $\mathrm{Isom}_{\mathbb{Z}}(\pi_{1}\partial M,\pi_{1}\partial N)/(\psi\sim-\psi)$. Indeed, under the ${}\mathbb{Q}\cup\{\infty\}$-parametrization, either $\overline{\psi}(s)=s$, or $\overline{\psi}(s)=-s$ for all $s\in\mathbb{Q}\cup\{\infty\}$.
In addition, $\overline{\psi}$ sends the rational null-homologous boundary slope of $M$ to the rational null-homologous boundary slope of $N$ by Lemma 4.4 (2). Since the linking number of the two components of $\mathcal{L}$ is $3$, the rational null-homologous boundary slope of $M$ corresponds to $\frac{3^{2}}{m-\frac{1}{2}}=\frac{18}{2m-1}$ in the ${}\mathbb{Q}\cup\{\infty\}$-parametrization. Similarly, the rational null-homologous boundary slope of $N$ corresponds to $\frac{3^{2}}{m-\frac{1}{2}}=\frac{18}{2n-1}$. Thus, $\overline{\psi}(\frac{18}{2m-1})=\frac{18}{2n-1}$. Recall that $\overline{\psi}$ is a multiplication by ±$\pm 1$ under the ${}\mathbb{Q}\cup\{\infty\}$-parametrization. Thus, $2m-1=\pm(2n-1)$. Note that $\mathcal{L}(r)\cong\mathcal{L}(-r)$ according to (6.16). Thus, regardless of the $\pm$-sign, $M=\mathcal{L}(\frac{2m-1}{2})\cong N=\mathcal{L}(\frac{2n-1}{2})$.
@@
We still consider the rational null-homologous boundary slopes. Let $r=\frac{p}{q}$ be a reduced fraction. Then (6.24) implies that $2q\alpha+2q\beta+(4q-p)\gamma$ is torsion in $H_{1}(L_{B}(0,r);\mathbb{Z})$. Thus, the slope represented by $\frac{4q-p}{q}=4-r$ on $\partial N(L_{B}^{(3)})$ is the rational null-homologous boundary slope on $\partial M$. Similarly, $4-s$ is the rational null-homologous boundary slope on $\partial N$. By Lemma 4.4 (2), $\overline{\psi}(4-r)=4-s$. Recall that either $\overline{\psi}(u)=u$, or $\overline{\psi}(u)=4-u$. Thus, either $s=r$ or $s=4-r$. Fortunately, [32, Lemma 2.2 (2)] proved that $L_{B}(0,r)\cong L_{B}(0,4-r)$. Thus, in both cases, $M\cong N$. ∎
@@
    $N$ is also irreducible and boundary-incompressible;
@@
As a standard application of the loop theorem, through attaching 2-handles to the compressible boundary components of $N$ and then capping off the emerging boundary spheres by 3-handles, we can obtain a boundary-incompressible 3-manifold $N^{+}$ such that $\pi_{1}N\cong\pi_{1}N^{+}$. Then $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N^{+}}$.
@@
Note that when $N$ is boundary-compressible, attaching a 2-handle along a compressible boundary curve will create an essential sphere, which makes $N^{+}$ reducible. Thus, $N$ is boundary-incompressible and $N=N^{+}$. Therefore, $N$ is also irreducible, which proves (1). (2) and (3) then follows from [57, Corollary 4.1]. ∎
@@
Let $M$ be a compact, orientable, irreducible 3-manifold with empty or incompressible toroidal boundary, which is not a closed $Sol$-manifold. Then there exists a minimal collection $\mathcal{T}$ of disjoint essential tori, such that each connected component of $M\setminus\mathcal{T}$ is either a Seifert fibered manifold or a finite-volume hyperbolic manifold. In addition, $\mathcal{T}$ is unique up to isotopy, and is called the JSJ-tori of $M$.
@@
The proof of [53, Theorem A] was stated for closed manifolds, while it applies to manifolds with boundary with no alteration.
Proposition A.3 ([56, Theorem 4.3] ).
Let $M$ and $N$ be compact, orientable, irreducible 3-manifolds with empty or incompressible toroidal boundary, which are not closed $Sol$-manifolds. Let $(\mathcal{G},\Gamma_{M})$ and $(\mathcal{H},\Gamma_{N})$ denote their JSJ-graphs of groups respectively. Suppose $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ is an isomorphism.
@@
    There is a graph isomorphism $f^{\dagger}:\Gamma_{M}\to\Gamma_{N}$, which preserves the hyperbolic/Seifert type of vertices.
@@
    ${\widehat{\mathcal{G}_{x}}}$${\widehat{\mathcal{H}_{f^{\dagger}(x)}}}$${\widehat{\pi_{1}M}}$${\widehat{\pi_{1}N}}$${\widehat{\pi_{1}N}}$$\scriptstyle{f_{x}}$$\scriptstyle{f}$$\scriptstyle{C_{h_{x}}}$
In fact, [56] was originally proven for closed manifolds, but the proof can be easily generalized to the bounded case, see [48, Theorem 5.3] or [57, Theorem 5.6].
@@
    $N$ is also a mixed 3-manifold.
@@
(1) By Lemma A.1, $N$ is irreducible and has empty or incompressible toral boundary. In addition, by [54, Theorem 8.4], $N$ is not a closed $Sol$-manifold. Therefore, $N$ satisfy the requirements for Proposition A.3. In particular, Proposition A.3 (1) implies that $N$ is also a mixed 3-manifold.
(2) We follow the notation in Proposition A.3. $(\mathcal{G},\Gamma_{M})$ and $(\mathcal{H},\Gamma_{N})$ denote the JSJ-graphs of groups, $f^{\dagger}:\Gamma_{M}\to\Gamma_{N}$ denotes the graph isomorphism, and $f_{x}:\widehat{\mathcal{G}_{x}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\mathcal{H}_{f^{\dagger}(x)}}$ denotes the induced isomorphism for each $x\in\Gamma_{M}$. For each vertex $v\in V(\Gamma_{M})$ corresponding to a hyperbolic piece, according to Proposition 2.2, $\widehat{\pi_{1}M_{v}}\cong\widehat{\pi_{1}N_{f^{\dagger}(v)}}$ implies that $M_{v}$ and $N_{f^{\dagger}(v)}$ have the same number of boundary components. Let $\partial^{hyp}M$ and $\partial^{hyp}N$ denote the boundary components belonging to hyperbolic pieces. Together with the graph isomorphism $f^{\dagger}:\Gamma_{M}\to\Gamma_{N}$,

### v3

@@ 6.4. The Berge manifold
Proposition 6.13 shows that there are three different slopes on one boundary component of $S^{3}\setminus\mathring{N}(L_{A})$ such that the corresponding Dehn filling is homeomorphic to $D^{2}\times S^{1}$. According to Theorem A, there exist three different slopes $c_{1},c_{2},c_{3}$ on one boundary component $\partial_{0}N$ of $N$ such that $\widehat{\pi_{1}N_{c_{1}}}\cong\widehat{\pi_{1}N_{c_{2}}}\cong\widehat{\pi_{1}N_{c_{3}}}\cong\widehat{\pi_{1}(D^{2}\times S^{1})}\cong\widehat{\mathbb{Z}}$. Note that the group $\mathbb{Z}$ is profinitely rigid among all finitely generated, residually finite groups. So according to Proposition 2.2, we derive that $\pi_{1}N_{c_{1}}\cong\pi_{1}N_{c_{2}}\cong\pi_{1}N_{c_{3}}\cong\mathbb{Z}$. For each $i=1,2,3$, $N_{c_{i}}$ is a compact orientable 3-manifold whose boundary is a torus, and the only such manifold with infinite cyclic fundamental group is $D^{2}\times S^{1}$. Thus, $N$ also admits three different Dehn fillings on $\partial_{0}N$ yielding $D^{2}\times S^{1}$, and Proposition 6.13 implies that $N\cong S^{3}\setminus\mathring{N}(L_{A})$. ∎
@@ 6.5. Detecting Klein bottles and Dehn surgeries on $\mathcal{L}$ and $L_{B}$
Recall the notations that $\mathcal{N}$ is the orientable $I$-bundle over Klein bottle, and $M_{1}$ is the complement of the Whitehead link $\mathcal{W}$ in $S^{3}$. In the following context, we denote $M_{3_{1}}$ as the complement of the trefoil in $S^{3}$, i.e. $M_{3_{1}}$ is a Seifert fibered space $(D^{2};(2,1),(3,1))$.
@@
(2) According to (6.15), $L_{B}(0,r,0)\cong L_{B}(0,0,r)$ and $L_{B}(0,r,4)\cong L_{B}(0,4,r)$. Thus, they can be viewed as Dehn fillings on $L_{B}(0,0)$ and $L_{B}(0,4)$. According to (1), $L_{B}(0,0,r)$ and $L_{B}(0,4,r)$ can be decomposed by an embedded torus into $\mathcal{N}\cup M^{\prime}$, where $M^{\prime}$ is a Dehn filling of $M_{1}$ on one of its boundary components $\partial_{1}M_{1}$. The Dehn fillings of $M_{1}$ on $\partial_{1}M_{1}$ have been classified by [37, Table A.1]. In fact, there is exactly one slope yielding the solid torus $D^{2}\times S^{1}$, while the other slopes yield either a hyperbolic manifold, a Seifert fibered space with hyperbolic base orbifold, or a graph manifold with non-trivial JSJ-decomposition. In other words, there is only one element $r_{0}\in\mathbb{Q}\cup\{\infty\}$ such that $L_{B}(0,0,r_{0})\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$, which is indeed non-aspherical; while for any $r\neq r_{0}\in\mathbb{Q}\cup\{\infty\}$, $L_{B}(0,0,r)\cong\mathcal{N}\cup_{T^{2}}M^{\prime}$ is aspherical. We claim that $r_{0}=\infty$. In fact, by (6.15), $L_{B}(0,0,\infty)\cong L_{B}(0,\infty,0)$ is a Dehn filling on $L_{B}(0,\infty)$. [32, Lemma 2.3 (1)] showed that $L_{B}(0,\infty)\cong\mathcal{N}$, and thus $L_{B}(0,\infty,0)\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$. Similarly, there is only one element $r_{4}\in\mathbb{Q}\cup\{\infty\}$ such that $L_{B}(0,4,r_{4})\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$ is spherical, while for any $r\neq r_{4}\in\mathbb{Q}\cup\{\infty\}$, $L_{B}(0,4,r_{4})\cong\mathcal{N}\cup_{T^{2}}M^{\prime}$ is aspherical. One can similarly show that $r_{4}=\infty$ by [32, Lemma 2.3 (1)]. Thus, this proves (2).
(3) follows from [37, Table 3]. In fact, the link $\mathcal{L}$ is obtained by a $\frac{1}{2}-$surgery on one component of the three-chain link (Figure 10) in $S^{3}$. Thus, for any $r,s\in\mathbb{Q}\cup\{\infty\}$, $\mathcal{L}(r,s)$ can be viewed as $(-\frac{1}{2},r-2,s-2)$-surgery on three components of the three-chain link, for which [37] presents a complete classification of exceptional Dehn surgeries.
@@
    Let $c_{1},c_{2}\in\text{slp}(T)$ and $c_{1}^{\prime},c_{2}^{\prime}\in\text{slp}(T^{\prime})$ such that $\Delta(c_{1},c_{2})=\Delta(c_{1}^{\prime},c_{2}^{\prime})=4$. Then there exist exactly two (equivalence classes of) isomorphisms $\psi_{1},\psi_{2}\in\mathrm{Isom}_{\mathbb{Z}}(\pi_{1}(T),\pi_{1}(T^{\prime}))/(\psi\sim-\psi)\approx\mathrm{PGL}(2,\mathbb{Z})$ such that $\overline{\psi}_{i}(\{c_{1},c_{2}\})=\{c_{1}^{\prime},c_{2}^{\prime}\}$.
@@
We can now list all possible elements in $E$. Suppose $c=[p\mu+q\lambda]\in E$, where $(p,q)$ is a pair of coprime integers. Without loss of generality, we may assume $p\geq 0$. Then $\Delta(c_{1},c)=\left|q\right|\leq 4$, and $\Delta(c_{2},c)=\left|4p-q\right|\leq 4$. It is easy to verify that $(p,q)\in\{(0,\pm 1),(1,0),(1,1),(1,2),(1,3),(1,4)\}$. Thus, $E\subseteq\{[\mu],[\mu+\lambda],[\mu+2\lambda],[\mu+3\lambda],[\mu+4\lambda],[\lambda]\}$, and it is easy to verify that $\mu]$ and $\mu+4\lambda]$ is the unique pair realizing geometric intersection number $4$.
@@
Therefore, $H_{1}(\mathcal{L}(n-\frac{1}{2});\mathbb{Z})\not\cong H_{1}(L_{B}(0,r);\mathbb{Z})$, which implies $\widehat{\pi_{1}\mathcal{L}(n-\frac{1}{2})}\not\cong\widehat{\pi_{1}L_{B}(0,r)}$ by Lemma 2.1.
Now we consider the family $\mathcal{L}(n-\frac{1}{2})$ ($n\in\mathbb{Z}$). Let $M=\mathcal{L}(m-\frac{1}{2})$ and $N=\mathcal{L}(n-\frac{1}{2})$, and suppose $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then according to Theorem A, there exists an isomorphism $\psi:\pi_{1}\partial M\to\pi_{1}\partial N$ such that $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ for all $c\in\text{slp}(\partial M)$. In particular, $\overline{\psi}(\mathcal{E}_{k}(M,\partial M))=\mathcal{E}_{k}(N,\partial N)$ according to Proposition 5.11. We follow the $\mathbb{Q}\cup\{\infty\}$-parametrization for $\text{slp}(\partial M)$ and $\text{slp}(\partial N)$ as above. Recall that $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{k}(M,\partial M)\}=\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{k}(N,\partial N)\}=4$ by Lemma 6.21. Thus, according to Corollary 6.19 (3) and Lemma 6.22 (1), $\{-2,2\}$ is the unique pair of elements in $\mathcal{E}_{k}(M,\partial M)$, as well as in $\mathcal{E}_{k}(N,\partial N)$, that realizes the maximal geometric intersection number $4$. Since $\overline{\psi}:\mathcal{E}_{k}(M,\partial M)\xrightarrow{1:1}\mathcal{E}_{k}(N,\partial N)$ preserves the geometric intersection number by Lemma 6.2, we derive that $\overline{\psi}(\{-2,2\})=\{-2,2\}$. According to Lemma 6.22 (2), there are only two possible candidates for $\psi$ in $\mathrm{Isom}_{\mathbb{Z}}(\pi_{1}\partial M,\pi_{1}\partial N)/(\psi\sim-\psi)$. Indeed, under the $\mathbb{Q}\cup\{\infty\}$-parametrization, either $\overline{\psi}(s)=s$, or $\overline{\psi}(s)=-s$ for all $s\in\mathbb{Q}\cup\{\infty\}$.
In addition, $\overline{\psi}$ sends the rational null-homologous boundary slope of $M$ to the rational null-homologous boundary slope of $N$ by Lemma 4.4 (2). Since the linking number of the two components of $\mathcal{L}$ is $3$, the rational null-homologous boundary slope of $M$ corresponds to $\frac{3^{2}}{m-\frac{1}{2}}=\frac{18}{2m-1}$ in the $\mathbb{Q}\cup\{\infty\}$-parametrization. Similarly, the rational null-homologous boundary slope of $N$ corresponds to $\frac{3^{2}}{m-\frac{1}{2}}=\frac{18}{2n-1}$. Thus, $\overline{\psi}(\frac{18}{2m-1})=\frac{18}{2n-1}$. Recall that $\overline{\psi}$ is a multiplication by $\pm 1$ under the $\mathbb{Q}\cup\{\infty\}$-parametrization. Thus, $2m-1=\pm(2n-1)$. Note that $\mathcal{L}(r)\cong\mathcal{L}(-r)$ according to (6.16). Thus, regardless of the $\pm$-sign, $M=\mathcal{L}(\frac{2m-1}{2})\cong N=\mathcal{L}(\frac{2n-1}{2})$.
@@
We still consider the rational null-homologous boundary slopes. Let $r=\frac{p}{q}$ be a reduced fraction. Then, (6.24) implies that $2q\alpha+2q\beta+(4q-p)\gamma$ is torsion in $H_{1}(L_{B}(0,r);\mathbb{Z})$. Thus, the slope represented by $\frac{4q-p}{q}=4-r$ on $\partial N(L_{B}^{(3)})$ is the rational null-homologous boundary slope on $\partial M$. Similarly, 4s4-s is the rational null-homologous boundary slope on $\partial N$. By Lemma 4.4 (2), $\overline{\psi}(4-r)=4-s$. Recall that either $\overline{\psi}(u)=u$, or $\overline{\psi}(u)=4-u$. Thus, either $s=r$ or s=4rs=4-r. Fortunately, [32, Lemma 2.2 (2)] proved that $L_{B}(0,r)\cong L_{B}(0,4-r)$. Thus, in both cases, $M\cong N$. ∎
@@
    NN is also irreducible and boundary-incompressible;
@@
As a standard application of the loop theorem, through attaching 2-handles to the compressible boundary components of $N$ and then capping off the emerging boundary spheres by 3-handles, we can obtain a boundary-incompressible 3-manifold $N^{+}$ such that $\pi_{1}N\cong\pi_{1}N^{+}$. Then, $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N^{+}}$.
@@
Note that when $N$ is boundary-compressible, attaching a 2-handle along a compressible boundary curve will create an essential sphere, which makes $N^{+}$ reducible. Thus, $N$ is boundary-incompressible and $N=N^{+}$. Therefore, $N$ is also irreducible, which proves (1). (2) and (3) then follow from [56, Corollary 4.2]. ∎
@@
Let $M$ be a compact, orientable, irreducible 3-manifold with empty or incompressible toroidal boundary. Then there exists a minimal collection $\mathcal{T}$ of disjoint essential tori, such that each connected component of $M\setminus\mathcal{T}$ is either a Seifert fibered manifold or a finite-volume hyperbolic manifold. In addition, $\mathcal{T}$ is unique up to isotopy, and is called the JSJ-tori of $M$.
@@
The proof of [52, Theorem A] was stated for closed manifolds, while it applies to manifolds with toral boundary with no alteration.
Recall that for any group $G$ and any element $g\in G$, we use $C_{g}(x)=gxg^{-1}$ to denote the conjugation.
Proposition A.3 ([55, Theorem 4.3]).
Let $M$ and $N$ be compact, orientable, irreducible 3-manifolds with empty or incompressible toral boundary, which are not closed $Sol$-manifolds. Let $(\mathcal{G},\Gamma_{M})$ and $(\mathcal{H},\Gamma_{N})$ denote their JSJ-graphs of groups respectively. Suppose $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ is an isomorphism.
@@
    There is a graph isomorphism $f^{\dagger}:\Gamma_{M}\to\Gamma_{N}$, which preserves the hyperbolic/Seifert type of the vertices.
@@
    $^{\widehat{\mathcal{G}_{x}}}$$^{\widehat{\mathcal{H}_{f^{\dagger}(x)}}}$$^{\widehat{\pi_{1}M}}$$^{\widehat{\pi_{1}N}}$$^{\widehat{\pi_{1}N}}$$\scriptstyle{f_{x}}$$\scriptstyle{f}$$\scriptstyle{C_{h_{x}}}$
  4. (4)
    ([56, Theorem 5.5]) For any $e\in E(\Gamma_{M})$ whose endpoints are denoted by $u$ and $v$, there exist $g\in\widehat{\mathcal{H}_{f^{\dagger}(u)}}$ and $h\in\widehat{\mathcal{H}_{f^{\dagger}(v)}}$ such that the following diagram commutes.
    $^{\widehat{\mathcal{G}_{u}}}$$\widehat{\mathcal{G}_{e}}\cong\widehat{\mathbb{Z}}^{2}$$^{\widehat{\mathcal{G}_{v}}}$$^{\widehat{\mathcal{H}_{f^{\dagger}(u)}}}$$\widehat{\mathcal{H}_{f^{\dagger}(e)}}\cong\widehat{\mathbb{Z}}^{2}$$^{\widehat{\mathcal{H}_{f^{\dagger}(v)}}}$$\scriptstyle{C_{g}\circ f_{u}}$$\scriptstyle{f_{e}}$$\scriptstyle{C_{h}\circ f_{v}}$
In fact, [55] was originally proven for closed manifolds, but the proof can be easily generalized to the bounded case, see also [48, Theorem 5.3].
@@
    NN is also a mixed 3-manifold.
@@
(1) By Lemma A.1, $N$ is irreducible and has empty or incompressible toral boundary. In addition, by [53, Theorem 8.4], $N$ is not a closed $Sol$-manifold. Therefore, $N$ satisfy the requirements for Proposition A.3. In particular, Proposition A.3 (1) implies that $N$ is also a mixed 3-manifold.
(2) We follow the notation in Proposition A.3: $(\mathcal{G},\Gamma_{M})$ and $(\mathcal{H},\Gamma_{N})$ denote the JSJ-graphs of groups, $f^{\dagger}:\Gamma_{M}\to\Gamma_{N}$ denotes the graph isomorphism, and $f_{x}:\widehat{\mathcal{G}_{x}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\mathcal{H}_{f^{\dagger}(x)}}$ denotes the induced isomorphism for each $x\in\Gamma_{M}$. For each vertex $v\in V(\Gamma_{M})$ corresponding to a hyperbolic piece, according to Proposition 2.3, $\widehat{\pi_{1}M_{v}}\cong\widehat{\pi_{1}N_{f^{\dagger}(v)}}$ implies that $M_{v}$ and $N_{f^{\dagger}(v)}$ have the same number of boundary components. Let $\partial^{hyp}M$ and $\partial^{hyp}N$ denote the boundary components belonging to hyperbolic pieces. Together with the graph isomorphism $f^{\dagger}:\Gamma_{M}\to\Gamma_{N}$,